Marst Corporation's budgeted production in units and budgeted raw materials purchases over the next three months are given below:

January February March
Budgeted production (in units) 54,000 ? 80,000
Budgeted raw materials purchases (in pounds) 115,800 141,800 197,800


Two pounds of raw materials are required to produce one unit of product. The company wants raw materials on hand at the end of each month equal to 30% of the following month's production needs. The company is expected to have 31,000 pounds of raw materials on hand on January 1. Budgeted production for February should be:

rev: 10_2
Answer #1
Let the Budgeted production for February be X
141800 = 2X+(80000*2*30%)-0.6X
141800 = 2X+48000-0.6X
2X-0.6X = 141800-48000
1.4X = 93800
X = 93800/1.4 = 67000
Budgeted production for February should be 67,000 units